You have bought 500 shares of XYZ Corp. at $36.71 per share on 70% margin. The maintenance margin is 40%. Next day, the price of the stock drops to $17.30 per share. if a margin call is triggered, calculate the value of the deposit required to service the margin call. If no margin call is issued, enter $0.00 as your answer.

Thank you!
Answer #1
A B C=A*B D=70%*C E=18355-12849 F=C-E G=40%*C
Day Share Price Number of Shares Account Value Initial Margin Margin Loan Account Equity Maintenance Margin Required
0 $36.71 500 $18,355 $12,849 $5,507 $12,849 $7,342
1 $17.30 500 $8,650 $5,507 $3,144 $3,460
Maintenance Margin required is more than Account Equity .
Hence there will be margin Call
Value of deposit required to service margin call $317 (3460-3144)
Margin Call will be when :
Account Value =Margin Loan/(1-MMR)
MMR=Minimum maintenance Margin=0.4
Account Value =5507/(1-0.4) $9,177.50
Share Price =9177.50/500= $18.36